OpenStudy (anonymous):

State whether the given measurements determine zero, one, or two triangles. C = 30°, a = 32, c = 16

OpenStudy (anonymous):

sin(A)/a=sin(C)/c sin(A)=a*sin(C)/c

OpenStudy (anonymous):

did you understand?

OpenStudy (anonymous):

I think so. I got an answer of 1? Does that mean one triangle? :P

OpenStudy (anonymous):

you got sin(A)=1 so A=90 degrees or 90+360 or 90+2*360... but in triangles the angles must be smaller than 90. so there is 1 possible value of A. therefore, 1 possible triangle

OpenStudy (anonymous):

I actually think I understand that. Thank you man. I really appreciate your help.

OpenStudy (anonymous):

you are welcome